# Understanding while Exploring: Semantics-driven Active Mapping

### v1

@@ 3.2 Exploration Planning
To enable efficient semantic reconstruction, we design an exploration planning module that actively selects informative viewpoints. Each candidate pose is evaluated using two criteria: geometric coverage, measured by silhouette completeness, and semantic uncertainty, quantified by entropy. These criteria approximate information gain [10, 12], which measures the expected reduction in uncertainty from new observations. While computing true information gain is intractable in high-dimensional semantic maps [96], our entropy- and coverage-based approximations allow efficient real-time scoring of candidate viewpoints. To keep computation efficient, we maintain a dynamic candidate pool and adopt a coarse-to-fine sampling strategy that first explores broadly, then refines with denser sampling. We now detail the geometric and semantic exploration criteria, the overall scoring formulation, and the implementation of candidate management.
@@
We adopt ActiveGAMER’s [74] exploration criterion formulation to evaluate the geometric coverage of candidate viewpoints. Given a candidate viewpoint $v$, we compute its exploration criterion $\mathcal{I}^{v}_{geo}$ based on the rendered silhouette $S^{v}$ with respect to the up-to-date Semantic Gaussian Map. The number of missing pixels in the rendered silhouette, denoted as $N_{S^{v}}$, quantifies the exploration criterion for the candidate viewpoint, which is formulated as:
@@
$\mathcal{I}^{v}_{\text{seman}}=\sigma\left(\sum\nolimits_{\mathbf{p}}H^{v}(\mathbf{p}\right),$ (7)
@@ 4.1 Experimental Setup
We use Habitat [97] to generate RGB-D frames and OneFormer [17] for semantic segmentation. Frames are captured at $680\times 1200$ resolution with $60^{\circ}$ vertical and $90^{\circ}$ horizontal FOV. The Exploration Map uses a voxel size of 5 cm.
We evaluate on three photorealistic datasets: Replica [98], ReplicaSLAM, and MP3D [99]. Replica includes high-fidelity meshes and 101 semantic classes; we use 8 scenes from [44]. ReplicaSLAM provides predefined camera trajectories for the same 8 scenes. MP3D includes 40 semantic classes; we use 5 scenes for 3D reconstruction evaluation. Each experiment runs for 2,000 steps on Replica and 5,000 on MP3D, with early termination if the exploration candidate pool is exhausted.
@@
We follow SGS-SLAM’s [7] evaluation protocol and compute the Average Mean Intersection over Union (mIoU) by mapping the rendered semantic predictions to ground-truth categories within each test view. We also evaluate per-pixel semantic classification using Top-1 and Top-3 Accuracy, and assess the complete category distribution (not restricted to categories present in the given image) using Mean Average Precision (mAP) and F-1 score.
@@
We evaluate geometric reconstruction using three metrics: Accuracy (cm), Completion (cm), and Completion ratio (%) with a 5 cm threshold. These are computed by uniformly sampling 3D points from both the ground-truth mesh and the reconstructed Gaussian Map. For measuring rendering quality, we use PSNR, SSIM, LPIPS and Depth L1 (D-L1).
To the best of our knowledge, this is the first work to address dense active semantic mapping using 3D Gaussian Splatting. We evaluate the effectiveness of our system against two categories of baselines: (1) semantic SLAM methods based on NeRF or 3DGS, which focus on segmentation and rendering quality; and (2) geometry-based active mapping methods, which prioritize 3D reconstruction accuracy. We conducted all the experiments on 2 NVIDIA RTX A6000 GPUs. Additional implementation details and results are provided in the supplementary material.
@@ 4.2 Semantic Segmentation Evaluation
NIDS-SLAM [83] ReplicaSLAM GT Train 2000 82.37
DNS-SLAM [86] ReplicaSLAM GT Train 2000 84.77
SNI-SLAM [85] ReplicaSLAM GT Train 2000 87.41
SGS-SLAM [7] ReplicaSLAM GT Train 2000 92.72
OneFormer [17] ReplicaSLAM GT Novel 3000 65.41
@@
SGS-SLAM [7] Replica Pred. Novel 2000 80.42 18.70
@@
SSMI [100] MP3D GT Train 36.14
TARE [101] MP3D GT Train 31.70
Zhang et al. [92] MP3D GT Train 42.92
@@
ReplicaSLAM. We evaluate on 4 scenes following the SGS-SLAM protocol [7], which compares rendered semantic masks to ground-truth labels visible in each view (Table 1 yellow). Our setup differs from baselines in three key ways: (1) we use pseudo labels from OneFormer [17] instead of ground-truth; (2) we evaluate on views unseen during training; and (3) we train using only one-third of the images. Despite these constraints, our method performs comparably to fully supervised baselines by fusing noisy predictions across views into a coherent semantic map.
Replica (Novel Views). To assess generalization, we generate new trajectories near the SLAM trajectories, following the instructions of SplaTAM. Table 1 blue compares three settings: (1) SGS-SLAM retrained with OneFormer labels; (2) our method without active exploration, showing the benefits of sparse semantic representation; and (3) our full pipeline with active exploration, which achieves better segmentation with fewer steps. Our method consistently outperforms the baseline, and Figure 2 shows improved alignment, density, and boundary quality in a Replica scene, office0.
@@
As shown in Figure 3, SGS-SLAM and similar methods use color encoding to represent semantic labels, which often blend during multi-view fusion and introduce arbitrary (yellow boxes). To recover labels, they apply nearest-color matching using either local label conversion, which maps to the nearest color among ground-truth classes in the current view, or global label conversion, which considers all ground-truth classes in the scene. However, assuming access to view-specific ground-truth labels is unrealistic. Inconsistencies between local and global conversion are shown in cyan boxes.
MP3D.
@@
We also evaluate on 5 large indoor scenes from MP3D (Table 1 red). We do not know which scenes were used in [92], but we evaluate on a common set of labels, and report Average IoU. Table 1 shows active semantic mapping baselines [100, 101] from [92]. All baselines use ground truth labels during optimization. Despite relying on predicted labels and novel views, our method significantly outperforms all baselines. Figure 4 shows that our system produces clean and consistent semantic maps across complex indoor scenes.
3D Reconstruction and Novel View Synthesis.
We evaluate ActiveSGM 3D reconstruction and novel view synthesis on MP3D and Replica. On MP3D, our method achieves 1.56 cm accuracy and 97.35% completeness, surpassing ActiveGAMER [74] (1.66 cm, 95.32%). In novel view synthesis on Replica, ActiveSGM achieves an SSIM of 0.96, closely matching ActiveGAMER’s 0.97 despite not using a photometric refinement stage. This highlights ActiveSGM’s ability to maintain a balance between photometric quality and geometric fidelity. Full quantitative and qualitative results are provided in the supplement.
@@ 4.3 Ablation Studies
We perform ablation studies on two key components of our proposed method that influence semantic mapping performance: (1) the number of categories used in the sparse semantic representation, and (2) the effect of individual loss terms in optimizing semantic features. Experiments are conducted on the office0 and room0 scenes from the Replica dataset. As shown in the supplement, using more categories improves accuracy: top-$[5,8,16]$ yields $[83.06\%,83.34\%,84.08\%]$ in Average mIoU. We retain the top-16 categories for the best overall performance. For the loss, removing either Hellinger distance or cosine similarity reduces the Average mIoU to 82.26% and 82.70%, respectively. Using both terms together, accuracy reaches $84.08\%$, confirming their effectiveness.
@@ 5 Limitations and Conclusion
We presented ActiveSGM, the first dense active semantic mapping system built on a 3D Gaussian Splatting backbone. Our approach unifies geometry, appearance, and semantics, enabling efficient exploration and high-quality mapping with fewer observations. ActiveSGM addresses key challenges in active semantic mapping: it improves semantic coverage via semantic-uncertainty-guided exploration, reduces memory with top-$k$ sparse representations, and handles noisy predictions without ground-truth labels.
Despite its strong performance, several limitations remain. The system relies on pseudo labels from a pretrained model (OneFormer), which may introduce domain-specific errors. Semantic initialization from initial observations can be unreliable in occluded regions. Additionally, to stabilize training, we block gradients from the semantic loss to the geometry properties, limiting joint optimization. Future work will explore more balanced multi-task learning and adaptive semantic refinement.
Our code will be released upon acceptance. We hope ActiveSGM will serve as a foundation for future research in active mapping, semantic understanding, and real-world robot autonomy.
Supplementary Materials
In this supplementary document, we provide a detailed outline structured as follows: Section S.1 offers additional implementation details of ActiveSGM. Section S.2 summarizes additional assumptions made in this paper. Section S.3 presents comprehensive ablation studies on the key components influencing semantic segmentation performance. Section S.4 includes extended quantitative and qualitative results, along with a runtime analysis. Section S.5 provides licenses for the existing assets.
@@
Hardware and Software.  We conducted the experiments on a server with 2 NVIDIA RTX A6000 GPUs and an Intel i9-10900X CPU with 20 cores. Our ActiveSGM is implemented with python 3.8 and CUDA 11.7. Please refer to Section S.5 for more information about baselines and other packages we used. Our code will be released upon acceptance.
@@
Replica [98] Train 97.31 98.75 98.67 99.17 97.35 96.45 98.13 98.83 91.15
@@
OneFormer Finetuning Details.  Following the instructions from ActiveGAMER [74], we implemented the geometry-based exploration criterion to construct our fine-tuning dataset. Beginning from a random position, the agent performs 500 exploration steps, collecting 500 RGB-Semantic frame pairs per scene. We then fine-tuned OneFormer [17] separately on the collected data from Replica and MP3D, training for 3,000 steps per scene. The Replica dataset has 101 classes, while MP3D has 40. The fine-tuning process follows the official OneFormer tutorial provided by Hugging Face (https://huggingface.co/docs/transformers/main/en/model_doc/oneformer). The novel trajectories described in Table 1 of the main paper are used as the test set. These trajectories are distinct from those used for fine-tuning. The train/test Top-1 accuracy is reported in Table S.1.
@@
Composite with alpha blending # sparse mode has fewer iters.
@@
Local Path Planner.  We employed the Efficient Rapid-exploration Random Tree (RRT) proposed by NARUTO [70] for local path planing. Once the goal location is determined, we use an efficient RRT-based planner to find a path from the current state $s_{t}$ to the goal $s_{g}$, using the Exploration Map to measure collision and reachability. (Specifically, the agent should only move within the free voxels defined by the Exploration Map. Additionally, we enforce a collision buffer of 20 cm, ensuring the agent avoids regions that are too close to surrounding surfaces.) To speed up planning in large-scale 3D environments, we enhance standard RRT by also attempting direct connections between samples and the goal. This greatly improves efficiency.
Appendix S.2 Assumptions
Due to space constraints in the main paper, we summarize below some additional assumptions made in this work:
  • Perfect Localization: Since this work focuses on active mapping, we assume the robot’s pose is known throughout the process. In real-world deployments, a separate localization or tracking module would be required.
  • Perfect Execution: We assume the robot can perfectly follow the planned trajectory to reach the selected next-best-view. In practice, navigation errors should be considered for deployment.
  • Semantic Segmentation Model: Our system depends on an external semantic segmentation model (OneFormer in our case) to generate semantic predictions. If a stronger model is used, the performance of our pipeline can improve. Conversely, if the model has limited accuracy or generalization ability, it may degrade the quality of the semantic map. Fine-tuning is recommended for adapting to new domains.
Appendix S.3 Ablation Studies.
We present the full ablation studies in Table S.2, consistent with the discussion in Section 4.3 of the main paper. We also compare KL divergence and Hellinger distance, finding that the Hellinger distance is a more effective choice. Notably, KL divergence can lead to gradient vanishing due to the instability of the logarithmic function, necessitating gradient norm clipping during training.
Table S.2: Ablation of Semantic Components. Experiments on office0 and room0 from Replica to evaluate the impact of the number of retained categories (Top-k) and the use of Hellinger distance (H.D.), KL-Divergence (KL.) and cosine similarity (Cos.) in the semantic loss.
Top-5 83.06 95.66 99.68 94.79 74.24
Top-8 83.34 95.70 99.66 95.05 74.23
Top-16 84.08 95.68 99.73 94.92 74.73
Top-16 82.26 95.57 99.61 94.21 74.10
Top-16 82.70 95.62 99.73 94.40 72.43
Top-16 82.22 95.63 99.70 94.66 73.93
Top-16 84.08 95.68 99.73 94.92 74.73
@@
In this section, we present additional qualitative and quantitative results on both datasets.
S.4.1 Quantitative Results
Semantic Segmentation on ReplicaSLAM  We evaluate on 4 scenes following the SGS-SLAM protocol [7], which compares rendered semantic masks to ground-truth labels visible in each view. The full results are shown in the Table S.3, and have been summarized in Table 1 yellow.
Table S.3: Semantic Segmentation on ReplicaSLAM. Rendered semantics are evaluated on 4 scenes using the SGS-SLAM [7] protocol, which compares predictions to ground-truth categories visible per view. Our method uses semantic predictions from OneFormer and is evaluated on novel views.
@@
NIDS-SLAM [83] GT Train 2000 82.37 82.45 84.08 76.99 85.94
DNS-SLAM [86] GT Train 2000 84.77 88.32 84.90 81.20 84.66
SNI-SLAM [85] GT Train 2000 87.41 88.42 87.43 86.16 87.63
SGS-SLAM [7] GT Train 2000 92.72 92.95 92.91 92.10 92.90
OneFormer [17] GT Novel 3000 65.41 69.06 65.71 67.01 59.85
@@
Semantic Segmentation on Replica (Novel View)  To assess generalization, we generate new trajectories near the SLAM trajectories, following the instructions of SplaTAM [51]. We present the complete results in Table S.4, as a supplement to Table 1 blue in the main paper.
@@
OneFormer [17] Steps $\downarrow$ - - - - - - - - -
@@
SGS-SLAM [7] Steps $\downarrow$ 2000 2000 2000 2000 2000 2000 2000 2000 2000
mIoU (%$\uparrow$ 80.42 77.60 75.68 78.70 78.10 89.96 83.23 83.97 76.12
mAP (%$\uparrow$ 89.94 86.37 84.67 88.63 90.98 96.22 92.10 93.80 86.73
F-1 (%$\uparrow$ 18.70 18.35 15.06 19.03 17.68 18.02 25.28 18.47 17.69
@@
SSMI [100] GT Train 36.14 46.02 41.01 25.13 39.30 36.12 29.25 -
TARE [101] GT Train 31.70 42.01 36.86 23.86 32.51 31.70 23.27 -
Zhang et al. [92] GT Train 42.92 50.73 45.26 43.91 40.42 39.18 37.99 -
@@
NARUTO[70] Replica 1.61 1.66 97.20
ActiveGAMER [74] Replica 1.16 1.56 96.50
@@
FBE [58] MP3D / 9.78 71.18
UPEN [102] MP3D / 10.60 69.06
OccAnt [103] MP3D / 9.40 71.72
ANM [104] MP3D 7.80 9.11 73.15
NARUTO[70] MP3D 6.31 3.00 90.18
ActiveGAMER [74] MP3D 1.66 2.30 95.32
@@
SplaTAM [51] Replica 29.08 0.95 0.14 1.38
SGS-SLAM [7] Replica 27.14 0.94 0.16 7.09
NARUTO [70] Replica 26.01 0.89 0.41 9.54
ActiveGAMER [74] Replica 32.02 0.97 0.11 1.12
@@
NARUTO [70] MP3D 20.52 0.72 0.58 7.95
ActiveGAMER [74] MP3D 24.76 0.90 0.25 4.83
@@
Appendix S.5 Licenses for existing assets
@@
    Matterport3D Dataset [99]

### v2

@@ 3.2 Exploration Planning
To enable efficient semantic reconstruction, we design an exploration planning module that actively selects informative viewpoints. Each candidate pose is evaluated using two criteria: geometric coverage, measured by silhouette completeness, and semantic uncertainty, quantified by entropy. These criteria approximate information gain [cadena2016past, placed2023survey], which measures the expected reduction in uncertainty from new observations. While computing true information gain is intractable in high-dimensional semantic maps [charrow2015information], our entropy- and coverage-based approximations allow efficient real-time scoring of candidate viewpoints. To keep computation efficient, we maintain a dynamic candidate pool and adopt a coarse-to-fine sampling strategy that first explores broadly, then refines with denser sampling. We now detail the geometric and semantic exploration criteria, the overall scoring formulation, and the implementation of candidate management.
@@
We adopt ActiveGAMER’s [chen2025activegamer] exploration criterion formulation to evaluate the geometric coverage of candidate viewpoints. Given a candidate viewpoint $v$, we compute its exploration criterion $\mathcal{I}^{v}_{geo}$ based on the rendered silhouette $S^{v}$ with respect to the up-to-date Semantic Gaussian Map. The number of missing pixels in the rendered silhouette, denoted as $N_{S^{v}}$, quantifies the exploration criterion for the candidate viewpoint, which is formulated as:
@@
$\mathcal{I}^{v}_{\text{seman}}=\sigma\left(\sum\nolimits_{\mathbf{p}}H^{v}(\mathbf{p})\right),$ (7)
@@ 4.1 Experimental Setup
In this paper, we focus on evaluating the semantic segmentation accuracy of the proposed pipeline, particularly under noisy observation settings, which better reflect real-world conditions. Deploying and evaluating on physical robots is challenging due to the lack of well-annotated data—especially ground-truth semantic labels- for reliable evaluation. Consequently, prior methods are typically evaluated in simulation, making photorealistic environments the most practical and fair testbed for benchmarking. Following the common evaluation protocol, all of the following experiments are conducted in a simulated environment.
@@
We use Habitat [savva2019habitat] to generate RGB-D frames and OneFormer [jain2023oneformer] for semantic segmentation. Frames are captured at $680\times 1200$ resolution with $60^{\circ}$ vertical and $90^{\circ}$ horizontal FOV. The Exploration Map uses a voxel size of 5 cm.
We evaluate on three photorealistic datasets: Replica [straub2019replica], ReplicaSLAM, and MP3D [chang2017matterport3d]. Replica includes high-fidelity meshes and 101 semantic classes; we use 8 scenes from [sucar2021imap]. ReplicaSLAM provides predefined camera trajectories for the same 8 scenes. MP3D includes 40 semantic classes; we use 5 scenes for evaluation. Each experiment runs for 2,000 steps on Replica and 5,000 on MP3D, with early termination if the exploration candidate pool is exhausted.
@@
We follow SGS-SLAM’s [li2024sgs] evaluation protocol and compute the Average Mean Intersection over Union (mIoU) by mapping the rendered semantic predictions to ground-truth categories within each test view. In addition, we evaluate per-pixel semantic classification using Top-1 and Top-3 Accuracy, and assess the complete category distribution (not limited to the categories present in a given image) using Mean Average Precision (mAP) and F1-score. Owing to space constraints, we report only a subset of the results in Table 1; please refer to Section S.2 of the supplementary material for the full results.
@@
We evaluate geometric reconstruction using three metrics: Accuracy (cm), Completeness (cm), and Completeness ratio (%) with a 5 cm threshold. These are computed by uniformly sampling 3D points from both the ground-truth mesh and the reconstructed Gaussian Map. For measuring rendering quality, we use PSNR, SSIM, LPIPS and Depth L1 (D-L1).
Baselines. To the best of our knowledge, this is the first work to tackle dense active semantic mapping with 3D Gaussian Splatting (3DGS); direct, one-to-one baselines are therefore difficult to establish. We evaluate our system against three categories of baselines: (1) semantic SLAM methods based on NeRF or 3DGS, which primarily target segmentation and rendering quality; (2) a passive semantic mapping pipeline reconfigured from the strongest representative in (1), to highlight the benefits of our semantic representation and active exploration policy; and (3) geometry-based active mapping methods, which focus on maximizing 3D reconstruction accuracy.
All experiments were conducted on two NVIDIA RTX A6000 GPUs. Additional implementation details and extended results are provided in the supplementary material.
@@ 4.2 Semantic Segmentation Evaluation
NIDS-SLAM [haghighi2023nids] ReplicaSLAM GT Train 2000 82.37
DNS-SLAM [li2024dns] ReplicaSLAM GT Train 2000 84.77
SNI-SLAM [zhu2024sni] ReplicaSLAM GT Train 2000 87.41
SGS-SLAM [li2024sgs] ReplicaSLAM GT Train 2000 92.72
OneFormer [jain2023oneformer] ReplicaSLAM GT Novel 3000 65.41
@@
SGS-SLAM [li2024sgs] Replica Pred. Novel 2000 80.42 18.70
@@
SSMI [asgharivaskasi2023semantic] MP3D GT Train 36.14
TARE [cao2023representation] MP3D GT Train 31.70
Zhang et al. [zhang2024active] MP3D GT Train 42.92
@@
Comparison with NeRF/GS-based Semantic SLAM on ReplicaSLAM. Most existing NeRF/3DGS-based semantic mapping approaches—such as NIDS-SLAM[haghighi2023nids], DNS-SLAM[li2024dns], SNI-SLAM[zhu2024sni], and SGS-SLAM[li2024sgs]—are formulated as semantic SLAM systems, i.e., they jointly solve for localization and semantic mapping. We report results from SGS-SLAM [li2024sgs] only for reference and completeness since we assume perfect localization. We follow the SGS-SLAM evaluation protocol, which measures the agreement between rendered semantic masks and ground-truth labels visible in each view (Table 1 yellow). Our setup differs from prior baselines in three key aspects: (1) we use pseudo labels generated by OneFormer [jain2023oneformer] instead of ground truth; (2) we evaluate on views not seen during training; and (3) we train with only one-third of the available images. Despite these constraints, our method achieves performance comparable to fully supervised baselines by effectively fusing noisy predictions across views into a coherent semantic map.
Comparison with Passive Semantic Mapping on Replica (Novel Views). To enable a fairer comparison with semantic SLAM methods, we select SGS-SLAM as the strongest representative baseline and reconfigure it to function as a semantic mapping-only system: (1) We disable its tracking module and use ground truth poses. (2) Replace ground truth semantics with semantic predictions. This allows us to directly evaluate the semantic representation quality, which is a key contribution of our method. We then disable our own active exploration module to ensure a passive, mapping-only setting, enabling a fair apples-to-apples comparison. The results in Table 1 blue show that our full pipeline with active exploration achieves better segmentation with fewer observations/steps, demonstrating the effectiveness of our exploration strategy guided by semantic and geometric uncertainty.
@@
As shown in Figure 3, SGS-SLAM and similar methods use color encoding to represent semantic labels, which often blend during multi-view fusion and introduce arbitrary labels, leading to misclassification and ambiguity (yellow boxes). To recover labels, they apply nearest-color matching using either local label conversion, which maps to the nearest color among ground-truth classes in the current view, or global label conversion, which considers all ground-truth classes in the scene. However, assuming access to view-specific ground-truth labels is unrealistic. Inconsistencies between local and global conversion are shown in cyan boxes.
@@
Comparison with Active Semantic Mapping on MP3D. We also include comparisons to recent active semantic mapping baselines, which are more aligned with our task definition. These further validate the advantage of our sparse semantic representation and active policy. We evaluate on 5 large indoor scenes from MP3D (Table 1 red). Table 1 shows active semantic mapping baselines [asgharivaskasi2023semantic, cao2023representation] from [zhang2024active]. We do not know which scenes were used by the baselines, but we evaluate on a common set of labels, and report Average IoU. All baselines use ground truth labels during optimization. Despite relying on predicted labels and novel views, our method significantly outperforms all baselines. Figure 4 shows that our system produces clean and consistent semantic maps across complex indoor scenes.
Comparison on 3D Reconstruction and Novel View Synthesis. We evaluate ActiveSGM 3D reconstruction and novel view synthesis on MP3D and Replica. On MP3D, our method achieves 1.56 cm accuracy and 97.35% completeness, surpassing ActiveGAMER [chen2025activegamer] (1.66 cm, 95.32%). In novel view synthesis on Replica, ActiveSGM achieves an SSIM of 0.96, closely matching ActiveGAMER’s 0.97 despite not using a photometric refinement stage. This highlights ActiveSGM’s ability to maintain a balance between photometric quality and geometric fidelity. Full quantitative and qualitative results are provided in the supplement.
Table 2: Ablation of Semantic Components. Experiments on office0 and room0 from Replica to evaluate the impact of the number of retained categories (Top-k) and the use of Hellinger distance (H.D.), KL-Divergence (KL.) and cosine similarity (Cos.) in the semantic loss.
Top-5 83.06 95.66 99.68 94.79 74.24
Top-8 83.34 95.70 99.66 95.05 74.23
Top-16 84.08 95.68 99.73 94.92 74.73
Top-16 82.26 95.57 99.61 94.21 74.10
Top-16 82.70 95.62 99.73 94.40 72.43
Top-16 82.22 95.63 99.70 94.66 73.93
Top-16 84.08 95.68 99.73 94.92 74.73
@@ 4.3 Ablation Studies
We perform ablation studies on two key components of our proposed method that influence semantic mapping performance: (1) the number of categories used in the sparse semantic representation, and (2) the effect of individual loss terms in optimizing semantic features. Experiments are conducted on the office0 and room0 scenes from the Replica dataset. As shown in Table 2: (1) Using more categories improves accuracy, we retain the top-16 categories for the best overall performance. (2) Removing either the Hellinger distance or cosine similarity reduces the Average mIoU. Using both terms together, accuracy reaches $84.08\%$, confirming their effectiveness. (3) We also compare KL divergence and Hellinger distance, finding that the Hellinger distance is a more effective choice. Notably, KL divergence can lead to gradient vanishing due to the instability of the logarithmic function, necessitating gradient norm clipping during training.
@@ 5 Limitations and Conclusion
Despite the strong performance of ActiveSGM we show above, several limitations remain: Perfect Localization: We assume known robot poses throughout the process; in real deployments, a separate localization or tracking module would be required. Perfect Execution: The robot is assumed to precisely follow planned trajectories; navigation errors should be considered for deployment. Semantic Segmentation Model: Our system relies on an external segmentation model (OneFormer) for semantic predictions, which may introduce domain-specific errors. Stronger or fine-tuned models can improve results, while weaker ones may degrade semantic quality. Limited Joint Optimization: To stabilize training, we block gradients from the semantic loss to geometric properties, decoupling the optimization of geometry and semantic features.
In conclusion, We presented ActiveSGM, the first dense active semantic mapping system built on a 3D Gaussian Splatting backbone. By unifying geometry, appearance, and semantics, ActiveSGM enables efficient exploration and high-quality mapping with fewer observations. It improves semantic coverage through uncertainty-guided exploration, reduces memory via top-$k$ sparse representations, and handles noisy predictions without ground-truth labels. We hope ActiveSGM will inspire future research in active mapping, semantic understanding, and autonomous robotic exploration. Our code is avaiable at https://github.com/lly00412/ActiveSGM.git.
Acknowledgement.
This research has been supported in part by the National Science Foundation under award 2024653.
Supplement
In this supplement, we provide a detailed outline structured as follows: Section S.1 offers additional implementation details of ActiveSGM. Section S.2 includes extended quantitative and qualitative results, along with a runtime analysis. Section S.3 provides justifications for the checklist items.
@@ Acknowledgement.
Hardware and Software.  We conducted the experiments on a server with 2 NVIDIA RTX A6000 GPUs and an Intel i9-10900X CPU with 20 cores. Our ActiveSGM is implemented with python 3.8 and CUDA 11.7. Please refer to Section S.3.4 for more information about baselines and other packages we used. Our code can be found at https://github.com/lly00412/ActiveSGM.git.
@@
Replica [straub2019replica] Train 97.31 98.75 98.67 99.17 97.35 96.45 98.13 98.83 91.15
@@
OneFormer Finetuning Details.  Following the approach of ActiveGAMER [chen2025activegamer], we implemented the geometry-based exploration criterion to construct our fine-tuning dataset. Beginning from a random position, the agent performs 500 exploration steps, collecting 500 RGB-Semantic frame pairs per scene. We then fine-tuned OneFormer [jain2023oneformer] separately on the collected data from Replica and MP3D, training for 3,000 steps per scene. The Replica dataset has 101 classes, while MP3D has 40. The fine-tuning process follows the official OneFormer tutorial provided by Hugging Face (https://huggingface.co/docs/transformers/main/en/model_doc/oneformer). The novel trajectories described in Table 1 of the main paper are used as the test set. These trajectories are distinct from those used for fine-tuning. The train/test Top-1 accuracy is reported in Table S.1.
@@
Composite with alpha blending # sparse mode needs fewer iters.
@@
Local Path Planner.  We employed the Efficient Rapid-exploration Random Tree (RRT) proposed by NARUTO [feng2024naruto] for local path planing. Once the goal location is determined, we use an efficient RRT-based planner to find a path from the current state $s_{t}$ to the goal $s_{g}$, using the Exploration Map to measure collision and reachability. (Specifically, the agent should only move within the free voxels defined by the Exploration Map. Additionally, we enforce a collision buffer of 20 cm, ensuring the agent avoids regions that are too close to surrounding surfaces.) To speed up planning in large-scale 3D environments, we enhance standard RRT by also attempting direct connections between samples and the goal. This greatly improves efficiency.
@@
S.2.1 Quantitative Results
Semantic Segmentation on ReplicaSLAM  We evaluate on 4 scenes following the SGS-SLAM protocol [li2024sgs], which compares rendered semantic masks to ground-truth labels visible in each view. The full results are shown in the Table S.2, and have been summarized in Table 1 yellow.
Table S.2: Semantic Segmentation on ReplicaSLAM. Rendered semantics are evaluated on 4 scenes using the SGS-SLAM [li2024sgs] protocol, which compares predictions to ground-truth categories visible per view. Our method uses semantic predictions from OneFormer and is evaluated on novel views.
@@
NIDS-SLAM [haghighi2023nids] GT Train 2000 82.37 82.45 84.08 76.99 85.94
DNS-SLAM [li2024dns] GT Train 2000 84.77 88.32 84.90 81.20 84.66
SNI-SLAM [zhu2024sni] GT Train 2000 87.41 88.42 87.43 86.16 87.63
SGS-SLAM [li2024sgs] GT Train 2000 92.72 92.95 92.91 92.10 92.90
OneFormer [jain2023oneformer] GT Novel 3000 65.41 69.06 65.71 67.01 59.85
@@
Semantic Segmentation on Replica (Novel Views)  To assess generalization, we generate new trajectories near the SLAM trajectories, following the instructions of SplaTAM [keetha2024splatam]. We present the complete results in Table S.3, as a supplement to Table 1 blue in the main paper.
@@
OneFormer [jain2023oneformer] Steps $\downarrow$ - - - - - - - - -
@@
SGS-SLAM [li2024sgs] Steps $\downarrow$ 2000 2000 2000 2000 2000 2000 2000 2000 2000
[…]
@@
SSMI [asgharivaskasi2023semantic] GT Train 36.14 46.02 41.01 25.13 39.30 36.12 29.25 -
TARE [cao2023representation] GT Train 31.70 42.01 36.86 23.86 32.51 31.70 23.27 -
Zhang et al. [zhang2024active] GT Train 42.92 50.73 45.26 43.91 40.42 39.18 37.99 -
@@
NARUTO[feng2024naruto] Replica 1.61 1.66 97.20
ActiveGAMER [chen2025activegamer] Replica 1.16 1.56 96.50
@@
FBE [yamauchi1997frontier] MP3D / 9.78 71.18
UPEN [georgakis2022uncertainty] MP3D / 10.60 69.06
OccAnt [ramakrishnan2020occupancy] MP3D / 9.40 71.72
ANM [yan2023active] MP3D 7.80 9.11 73.15
NARUTO[feng2024naruto] MP3D 6.31 3.00 90.18
ActiveGAMER [chen2025activegamer] MP3D 1.66 2.30 95.32
@@
SplaTAM [keetha2024splatam] Replica 29.08 0.95 0.14 1.38
SGS-SLAM [li2024sgs] Replica 27.14 0.94 0.16 7.09
NARUTO [feng2024naruto] Replica 26.01 0.89 0.41 9.54
ActiveGAMER [chen2025activegamer] Replica 32.02 0.97 0.11 1.12
@@
NARUTO [feng2024naruto] MP3D 20.52 0.72 0.58 7.95
ActiveGAMER [chen2025activegamer] MP3D 24.76 0.90 0.25 4.83
@@
Appendix S.3 Assets Used and Reproducibility
S.3.1 Experimental result reproducibility
We provide sufficient implementation details in Section S.1 in the supplement to make results reproducible. We build upon open source software, and our code can be found at https://github.com/lly00412/ActiveSGM.git.
S.3.2 Experimental setting/details
Please refer to Section 4.1 in the main paper.
S.3.3 Experiments compute resources
Please see the Hardware paragraph in Section S.1 in the supplement, and also refer to Sec S.2.2 for runtime analysis.
S.3.4 Licenses for existing assets
@@
    Matterport3D Dataset [chang2017matterport3d]